OpenStudy (anonymous):

Find the term containing x^5 in the expansion of (2x+3) (x-2)^6

OpenStudy (anonymous):

@zepdrix ^-^

zepdrix (zepdrix):

Hmm I haven't seen this type of problem before :d I'll have to think about it a sec.

OpenStudy (anonymous):

okie no biggie

zepdrix (zepdrix):

I guess we could use Binomial Theorem to turn the second part into a sum, and then distribute that sum to the first set of brackets. maybe.

OpenStudy (anonymous):

alrighty. do you wanna write it all out or should i help (but i think i'd be really bad help tbh)

zepdrix (zepdrix):

\[\large\rm (2x+3)\color{orangered}{(x-2)^6}\quad=(2x+3)\color{orangered}{\sum_{k=0}^{6}\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^k}\] Haha I dunno, I don't wanna steal all of the fun XD but it's a little harder to explain if I can't write the stuff out.

OpenStudy (anonymous):

ooh i could tell you the answer

OpenStudy (anonymous):

and you could work towards getting it

OpenStudy (astrophysics):

shouldn't it be the other way xD

OpenStudy (anonymous):

\[8\left(\begin{matrix}6 \\ 2\end{matrix}\right) x^5 - 6\left(\begin{matrix}6 \\ 1\end{matrix}\right)x^5 = 84x^5\]

OpenStudy (anonymous):

ahahahah yea probably

OpenStudy (anonymous):

but the back of my book has the answers i just don't know how to get them ;-;

zepdrix (zepdrix):

Ok I think we're on the right track then :) Let's see how this goes.

zepdrix (zepdrix):

Distributing,\[\large\rm =2x \sum(stuff) +3\sum(stuff)\] Ok with this step?

OpenStudy (anonymous):

kinda lol

OpenStudy (anonymous):

i mean what's the "stuff" supposed to be?

zepdrix (zepdrix):

The stuff from the orange sum posted before.

OpenStudy (anonymous):

ohh ok right right

zepdrix (zepdrix):

\[\large\rm =2x \sum\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^k +3\sum(stuff)\]Let's bring the 2 and x into the first sum, and combine with their respective values,\[\large\rm =\sum\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k+1}(-2)^{k+1} +3\sum(stuff)\] Ok lemme replace the other "stuff" as well at this point.

zepdrix (zepdrix):

\[\large\rm =\sum\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k+1}(-2)^{k+1} +\sum 3\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^{k}\]

OpenStudy (anonymous):

im confused :/

zepdrix (zepdrix):

By which part? :o

OpenStudy (anonymous):

ok im looking up when you did \[2xSigma(stuff)\]

OpenStudy (anonymous):

or whatever it was. i just don't get that part. where's the 2x coming from? and what's the "stuff" specifically

OpenStudy (anonymous):

also i don't understand where you get the 3

OpenStudy (anonymous):

basically everywhere down from there, i don't get

zepdrix (zepdrix):

We started with Binomial Theorem:\[\large\rm (2x+3)\color{orangered}{(x-2)^6}\quad=(2x+3)\color{orangered}{\sum_{k=0}^{6}\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^k}\]We're writing it like this for convenience,\[\large\rm =(2x+3)\color{orangered}{\sum_{k=0}^{6}stuff}\]

OpenStudy (anonymous):

ohhok

OpenStudy (anonymous):

ok ok now i get it

zepdrix (zepdrix):

Distribute the sum to each term inside the brackets,\[\large\rm =2x\color{orangered}{\sum_{k=0}^{6}stuff}+3\color{orangered}{\sum_{k=0}^{6}stuff}\]

zepdrix (zepdrix):

To bring the 2 and x inside the sum, you need to apply exponent rule to each, in order to combine the bases I mean :)

OpenStudy (anonymous):

so if you wrote it out, it'd be \[2x * (x^{6-r} * (-2)^r)\]

OpenStudy (anonymous):

and then added to what 3 distributed to that stuff would be

zepdrix (zepdrix):

Oh I switched to k again didn't I? woops >.<

OpenStudy (anonymous):

haha its fine

zepdrix (zepdrix):

For your general term? Ya that sounds right, but with the "choose" bracket thingy also, ya?\[\large\rm =2x\left[\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^k\right]+3\left[\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^k\right]\]

zepdrix (zepdrix):

Actually, these k's correspond to separate summations,\[\large\rm =2x\left[\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{6-k}(-2)^k\right]+3\left[\left(\begin{matrix}6 \\ r\end{matrix}\right)x^{6-r}(-2)^r\right]\]So I'mma be a little sloppy and use r for the other one. We want to find the value k in the first term that gives us our x^5, and our value of r in the second term that gives us our x^5.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so would you set 6-r to 5?

OpenStudy (anonymous):

so it'd be \[6-r\]

OpenStudy (anonymous):

=5

zepdrix (zepdrix):

\[\large\rm =\left[2\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{7-k}(-2)^k\right]+\left[3\left(\begin{matrix}6 \\ r\end{matrix}\right)x^{6-r}(-2)^r\right]\]The first term is a little trickier, because we want to bring the 2 and x in. (I added 1 to the 6-k exponent, that's how he joined in). For the other term? Yes that looks right :) 6-r = 5

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

can you give me a minute?

zepdrix (zepdrix):

ya

zepdrix (zepdrix):

imma grab some foods real quick :d

OpenStudy (anonymous):

okok so then for the first one it'd be r=2 and then the second one would be 1, right? and ok haha

zepdrix (zepdrix):

k=2, r=1. Ya that sounds right :)

OpenStudy (anonymous):

ok but did you see the answer

OpenStudy (anonymous):

its kinda....different from what we got

zepdrix (zepdrix):

Is it? Hmm, looks the same to me.

OpenStudy (anonymous):

it's \[8 \left(\begin{matrix}6 \\ 2\end{matrix}\right) x^5 - 6 \left(\begin{matrix}6 \\ 1\end{matrix}\right)x^5 = 84x^5\]

OpenStudy (anonymous):

wait ok how did they get that from r=2, r=1

zepdrix (zepdrix):

\[\large\rm =\left[2\left(\begin{matrix}6 \\ k\end{matrix}\right)x^{7-k}(-2)^k\right]+\left[3\left(\begin{matrix}6 \\ r\end{matrix}\right)x^{6-r}(-2)^r\right]\]We determined that k=2, r=1.\[\large\rm =\left[2\left(\begin{matrix}6 \\ 2\end{matrix}\right)x^{7-2}(-2)^2\right]+\left[3\left(\begin{matrix}6 \\ 1\end{matrix}\right)x^{6-1}(-2)^1\right]\]

zepdrix (zepdrix):

And then we just have to simplify a bit further, ya?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ok then um you would add the 4 to the 2?

OpenStudy (anonymous):

but i don't get that either i mean how are they getting 8

zepdrix (zepdrix):

add? 0_o

OpenStudy (anonymous):

omg no it's multiplying

zepdrix (zepdrix):

:)

OpenStudy (anonymous):

ok you're literally a genius you should get paid for this

OpenStudy (anonymous):

my tutor explained it in such a complicated way and did pascal's triangle and everything but this was sooo nice and easy

zepdrix (zepdrix):

lolol XD aren't you sweet. i do tutoring in real life sometimes :)) too busy with school'n right now though

OpenStudy (anonymous):

haha np~ good, i hope you get paid lots lol. anyway, i gotta go study some more but i might pop up later tonight. thanks for all your help!

zepdrix (zepdrix):

Well you would need to use Pascal's Triangle to do the REST of the problem. Like do you understand how to simply the 6choose2 and 6choose1?

OpenStudy (anonymous):

oh umm

OpenStudy (anonymous):

what? i dont get the last part

zepdrix (zepdrix):

\[\left(\begin{matrix}6 \\ 1\end{matrix}\right)=6\]And the other one, I'm not sure, I'd have to go down the triangle to see what the next number in the row is, I think it's 15 or something.

zepdrix (zepdrix):

\[\left(\begin{matrix}6 \\ 2\end{matrix}\right)=15,\qquad\qquad maybe\]

zepdrix (zepdrix):

Triangle is used to get those values ^

OpenStudy (anonymous):

oh yea

zepdrix (zepdrix):

Or you can use this if you're comfortable with it:\[\large\rm \left(\begin{matrix}n \\ r\end{matrix}\right)=\frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

yea i know how to do that haha

zepdrix (zepdrix):

ok good :) happy studying!

zepdrix (zepdrix):

and Christmas!

OpenStudy (anonymous):

haha thanks you too!